Nevey:

A student needs to make a rectangular cardboard piece that is 12 inches long. The area of the cardboard piece should be between 60 square inches to 300 square inches. The function f(w) relates the area of the cardboard, in square inches, to the width w in inches. Which of the following shows a reasonable domain for f(w)? 2 ≤ w ≤ 5 5 ≤ w ≤ 25 25 ≤ w ≤ 60 60 ≤ w ≤ 300

HarleyQuinn112:

area= L x w 60= 12 x w w=60/12=5 300 = 12 x w w = 300/12 = 25 so 25 or B is the correct answer

Nevey:

okie ty Harley :)

HarleyQuinn112:

anytime :)